Jefferson & Sons is evaluating a project that will increase annual sales by $450,000 and annual costs by $270,000. The project will initially require $130,000 in fixed assets that will be depreciated straight-line to a zero book value over the 5 year life of the project. The applicable tax rate is 40 percent. The annual operating cash flow for this project is $___. Round it to a whole dollar.
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(130,000/5)=$26,000

Annual operating cash flow=(Sale-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(450,000-270,000)(1-0.4)+(0.4*26,000)

=$118400